# Privacy Guarantees in Posterior Sampling under Contamination

### v2

@@ 5 Examples
We have seen in the previous section that empirically the approximated $\hat{\epsilon}_{n},\hat{\delta}_{n}$ follows the decreasing trend as in Theorem 5. In this section, we give a few examples that Theorem 5 is applicable theoretically. The verification of Assumptions 1 to 3 across different models shares a few common procedures. For Assumption 4 to hold for regression models (with covariates), one needs to control the leverage of the covariates $\bm{W}$. It suffices to assume $\|\bm{W}\|_{\infty}\leq n^{\iota}$, for some $\iota\in(0,1)$ that depends on the choice of $\phi_{n}$. However, for simplicity, we will still assume $\|\bm{W}\|_{\infty}\leq 1$, as we did in the simulations in Section 4. With $\|\bm{W}\|_{\infty}\leq 1$, Assumption 4 can be trivially satisfied by correctly choosing a heavy-tailed contamination density $g$, and Assumption 5 is usually easy to check for linear models.
Due to the length, we will only present the Bayesian quantile regression case in this section. We show three other cases in Appendix C that is contained in the supplement (Hu et al., 2024).
5.1 Bayesian Quantile Regression
In quantile regression, one would like to estimate the $\tau$-th quantile, denoted $q_{\tau}(X_{i}|\bm{W}_{i})$ of the random variable $X_{i}$ conditioned on the covariates $\bm{W}_{i}$. Due to Koenker and Bassett Jr (1978), solving the quantile regression problem is equivalent to solving the following optimization problem:
$\min_{\theta}\sum_{i}Q_{\tau}(x_{i}-\theta^{\top}\bm{w}_{i}),\quad Q_{\tau}(u)=\frac{1}{2}(|u|+(2\tau-1)u)$
This is equivalent to maximizing the likelihood
$\max_{\theta}\prod_{i}f_{\tau}(x_{i};\bm{w}_{i},\theta)$
where the probability density function of $X$ given $\tau,\theta,\bm{w}$ is
$f^{\tau}_{\theta}(x|\bm{w})=\tau(1-\tau)\exp\left(-Q_{\tau}(x-\theta^{\top}\bm{w})\right)$
describes the asymmetric Laplace distribution. Please refer to Yu and Moyeed (2001) for more details on Bayesian quantile regression. Firstly, we check Assumption 1 by using Proposition 3. We showed that, by Lemma 27, for any $\theta,\vartheta\in\Omega_{n}:=\{\theta\in\Omega:\|\theta-\theta^{*}\|_{2}\leq R_{n}\}$, such that $\|\theta-\vartheta\|_{\infty}\leq r$,
$\left|\sqrt{f^{\tau}_{\theta}(x|\bm{w})}-\sqrt{f^{\tau}_{\vartheta}(x|\bm{w})}\right|\leq M_{\vartheta,r}(x,\bm{w})\|\theta-\vartheta\|_{\infty}$
where $\tau^{*}=\max\{\tau,1-\tau\}$, $M_{n}=R_{n}+\|\theta^{*}\|_{2}$ and
Then by Lemma 28, assuming $\|\bm{w}\|_{\infty}\leq 1$,
Thus Assumption 1 can be satisfied with the choice $\phi_{n}=n^{-1/q},R_{n}=n^{k}$ for some $q>4$ and $k>0$ to be chosen for Assumption 3. It is not hard to check that the likelihood function satisfies the condition in Lemma 7, when $\theta$ is close enough to $\theta^{*}$, i.e., within $L_{2}$-distance of some constant $c_{r}$. Thus $S_{0,\alpha}(t_{n})$ contains an $L_{2}$-neighbourhood around $\theta^{*}$ with its radius proportional to $t_{n}$ up to a constant depending on $\theta^{*}$ and $c_{r}$. Then Assumption 2 is satisfied by applying Lemma 8 and Remark 12. For Assumption 3, we consider for $R_{n}$ large and apply Remark 13,
and we need
$R_{n}^{2}=n^{2k}\geq 4c_{1}n\phi_{n}^{2}=4c_{1}n^{(q-2)/q}$
which can be satisfied by choosing $k>(q-2)/2q$, for instance, $q=4$, $k>\tfrac{1}{4}$.
Finally, Lemma 29 also shows that $\phi_{n}$ and $\psi_{n}$ in Proposition 4 can be chosen to diminish at the same rate, similar to the conclusions as in the previous examples. Thus, we can choose $\psi_{n}\propto n^{-1/q}$ and $p_{n}=n^{-1/2q}$ which gives
$\epsilon_{n}\sim n^{-1/2q},\qquad\delta_{n}\sim\exp(-\tfrac{1}{2}c_{1}n^{1-2/q})$
where $c_{1}$ is a fixed constant and $q>4$.
@@ 6 Discussion
Differential privacy mechanisms in statistical inference can be roughly divided into two categories: 1. Local models, where data are privatized before collection and inference is conducted on the perturbed dataset; 2. Central models, where data are collected and processed without noise but the result is privatized before release.
@@
The focus of the ”local model”, or ”local differential privacy” as defined in Erlingsson, Pihur and Korolova (2014), quantifies the level of differential privacy provided by the mechanism on ”datasets” with only one data point. Thus contamination can be done individually before the dataset is gathered, Huber’s contamination model is a popular approach to local differential privacy. Our work, however, considers differential privacy from a central model perspective despite using a local privatization mechanism. Notably, the local differential privacy level only depends on the contamination probability $p_{n}$, but $p_{n}\rightarrow 0$ as $n\rightarrow\infty$. From the local model perspective, our setup will be asymptotically non-private as $n\rightarrow\infty$. However, from the central model perspective, with a trusted curator, our approach can ensure full differential privacy with asymptotically zero contamination rate.
@@
Huber’s contamination model for robust estimation by contaminating the dataset with a heavy-tailed distribution also introduces differential privacy to posterior sampling similar to central models that directly inject Laplace/Gaussian noise into the output. By contaminating the dataset with a heavy-tailed distribution, we can lift the assumption on bounded observation space. We further extend the result, through a probabilistic argument, to lift the bound on the parameter space and upper-bound the rate of convergence for $\epsilon$ and $\delta$.
@@
Arguably, privacy through contamination is more robust since the parameter/observation spaces are allowed to be unbounded and the introduced noise is mostly unaffected by the change in the size of the space, as shown in the simulations. Through simulation, a practitioner can estimate the level of differential privacy without having to hand-compute all the constants in the theorem. Despite having a couple of non-trivial assumptions for the theorem, we show in Section 5, how the assumptions could be verified mathematically and how the convergence rate of $\epsilon$ and $\delta$ can be derived under a few common regression setups.
Our main result (Theorem 5) in principle holds uniformly for all data size $n$ as long as the assumptions are satisfied. However, one should perceive it as an asymptotic result since $n$ has to be large for some assumption to hold in general. Shall the constants be improved will extend the practicality of our results. In addition, our main result is built primarily on the contraction of the posterior distribution. It is plausible to extend our result to other interesting models like misspecified models, infinite-dimensional models, etc, when the compatible posterior contraction results are present.
The underlying assumption of our analysis is that i.i.d. samples are drawn exactly from the target distribution. Applying results from Minami et al. (2016) can extend the our bounds non-exact samplers if the algorithm has an established convergence bound. The analysis of Markov chain convergence is an interesting area of research on its own with extensive literature, e.g., Andrieu et al. (2022). Thus we would like to leave the relevant analysis as future work and assume a ”good” sampler is readily available.
@@
$\mathbb{P}^{*}_{\theta^{*}}\left(\sup_{\theta\in A_{n}^{c}\cap\Omega_{n}}\prod_{i=1}^{n}d(\bm{x}_{i},\theta)\geq\exp({-c_{1}n\phi_{n}^{2}})\right)\leq(1+L(\phi_{n}))(1+N(\phi_{n}))\exp({-c_{2}n\phi_{n}^{2}})$
@@
If in addition that $\exp\left(-c_{2}n\phi_{n}^{2}\right)\leq\frac{1}{2},$ then
@@
Setting $p^{\prime}=0$, and we have
@@
See Lemma 2 of Shen and Wasserman (2001), the proof holds for any fixed $p\in[0,1)$. ∎
@@
For $\alpha\in(0,1]$, $S_{p,\alpha}(t)=\{\rho_{p,\alpha}(\theta^{*},\theta)\leq t)\}$, then $S_{0,\alpha}(t)\subseteq S_{p,\alpha}(t)$.
@@
Then apply Lemma 11 with the above inequality to give the bound. ∎
@@
By Lemma 13 and 14, and recall $t_{n}=c_{1}\phi_{n}^{2}/8$ defined in Assumption 2,
@@
and from Lemma 15 that $\eta_{n}\leq e^{2(1-p_{n})T_{n}\psi_{n}/p_{n}}.$ Let $\delta_{n}=\tfrac{4}{c_{3}}e^{-\frac{1}{2}c_{1}n\phi_{n}^{2}}$, $\varepsilon_{n}=2(1-p_{n})T_{n}\psi_{n}/p_{n}$ and $\epsilon_{n}=\varepsilon_{n}+\log(\delta_{n}+\exp(\varepsilon)).$ Note that since $\bm{X}$ and $\bm{Z}$ has the same cardinality, we apply Lemma 14 again to $m_{n}(A_{n}^{c},\bm{Z})$, we have
$\mathbb{P}\Bigl{(}\mathbb{P}_{\pi_{n}}(S|\bm{X})\leq e^{\epsilon_{n}}\mathbb{P}_{\pi_{n}}(S|\bm{Z})+\delta_{n}\Bigr{)}\geq$  $\,1-\mathbb{P}\left(\frac{m_{n}(A_{n}^{c},\bm{X})}{m_{n}(\Omega,\bm{X})}\geq\frac{4}{c_{3}}e^{-\frac{1}{2}c_{1}n\phi_{n}^{2}}\right)$
$\geq$  $\,1-2e^{-nc_{1}\alpha\phi_{n}^{2}/8}-C(n,\phi_{n})e^{-c_{2}n\phi_{n}^{2}}-c_{4}e^{-c_{1}n\phi_{n}^{2}}$
$\geq$  $\,1-C^{\prime}(n,\phi_{n})e^{-nc_{5}\phi_{n}^{2}}$
where $c_{5}=\min\{c_{1}\alpha/8,c_{2}\}$ and $C^{\prime}(n,\phi_{n})=2+C(n,\phi_{n})+c_{4}$. Note by definition of $C(n,\phi_{n})$ from Lemma 14, $C^{\prime}(n,\phi_{n})=20(\log_{2}n)^{2}$. On the other hand, if $n$ is large enough, then $C^{\prime}(n,\phi_{n})=6+c_{4}$ is independent of $n$, and the proof is complete. ∎
@@
Let $H(\bm{x}):=\left[\nabla_{\theta}f(\bm{x};\theta^{*})\right]\left[\nabla_{\theta}f(\bm{x};\theta^{*})\right]^{\top}$ and let $[H(\bm{x})]_{i,j}$ denote its $(i,j)$-th entry. Consider the sets $A_{i,j}:=\{x:[H(\bm{x})]_{i,j}\geq 0\}$, we have
$0<\left|\frac{[H(\bm{x})]_{i,j}}{k_{p_{n}}(\bm{x};\theta^{*})f(\bm{x};\theta^{*})}\right|\leq\frac{[H(\bm{x})]_{i,j}}{(1-p_{n})f^{2}(\bm{x};\theta^{*})},\text{ for }x\in A_{i,j}$
@@
$0<\left|\frac{[H(\bm{x})]_{i,j}}{k_{p_{n}}(\bm{x};\theta^{*})f(\bm{x};\theta^{*})}\right|\leq-\frac{[H(\bm{x})]_{i,j}}{(1-p_{n})f^{2}(\bm{x};\theta^{*})},\text{ for }x\in A_{i,j}^{c}.$
Note that $A_{i,j}$ and $A_{i,j}^{c}$ are both measurable with
$[I(\theta^{*})]_{i,j}=\int_{A_{i,j}\cup A^{c}_{i,j}}\frac{[H(\bm{x})]_{i,j}}{f^{2}(\bm{x};\theta^{*})}f(\bm{x};\theta^{*})\mathtt{d}x.$
Thus, apply dominated convergence theorem on both $A_{i,j}$ and $A_{i,j}^{c}$,
$\lim_{n\rightarrow\infty}\left[I_{p_{n}}(\theta^{*})\right]_{i,j}$ $\,=\lim_{n\rightarrow\infty}(1-p_{n})^{2}\left(\int_{A_{i,j}}\frac{[H(\bm{x})]_{i,j}}{k_{p_{n}}(\bm{x};\theta^{*})f(\bm{x};\theta^{*})}f(\bm{x};\theta^{*})\mathtt{d}x\right.$
$\,\qquad\qquad\qquad\qquad\left.+\int_{A_{i,j}^{c}}\frac{[H(\bm{x})]_{i,j}}{k_{p_{n}}(\bm{x};\theta^{*})f(\bm{x};\theta^{*})}f(\bm{x};\theta^{*})\mathtt{d}x\right)$
$\,=\int_{A_{i,j}\cup A_{i,j}^{c}}\frac{[H(\bm{x})]_{i,j}}{f^{2}(\bm{x};\theta^{*})}f(\bm{x};\theta^{*})\mathtt{d}x=[I(\theta^{*})]_{i,j}.$
@@ B.1 Results in Main Text
$D(u,\Omega):=\max\{N\in\mathbb{N}:\exists\text{ a set of $N$ points that is $u$-separate under $\|\cdot\|_{\infty}$}\}.$
@@ B.1.1 Proof for Lemma 3
The first half of this proof is based on several results commonly seen in Empirical processes literature (with some variation), see, for instance, Section 2.8 of Van der Vaart and Wellner (1996).
@@
$N_{[]}(uM^{(n)}_{u},\mathcal{F}(\Omega_{n}))\leq C_{d}\text{Vol}(\tilde{\Omega}_{n})u^{-d}$ (25)
@@
Let $\{\theta_{1},\dots,\theta_{k}\}$ be an $u$-cover of $\Omega_{n}$ with respect to $\|\cdot\|_{\infty}$. Consider the following functions, for $i=1,\dots,k$,
@@
Thus the set of $[L_{i},U_{i}]$ form a $uM^{(n)}_{u}$-bracketing for $\mathcal{F}(\Omega_{n})$ under the Hellinger distance $h(\cdot,\cdot)$, hence,
@@
To complete (25), we recall that $N(u,\Omega_{n})\leq D(u,\Omega_{n})$ always holds. If $\{\theta_{1},\dots,\theta_{k}\}$, $k:=D(u,\Omega_{n})$, is a set of $u$-separated points, then the balls of radius $u/2$ centred at $\{\theta_{1},\dots,\theta_{k}\}$ are disjoint and contained in $\tilde{\Omega}_{n}:=\{\theta\in\Omega:\exists\theta^{\prime}\in\Omega_{n},\|\theta-\theta^{\prime}\|_{2}\leq u/2\}$. Thus $D(u,\Omega_{n})$ is upper bounded by the ratio between the volume of the space $\tilde{\Omega}_{n}$ and the $k$-collection ball of radius $u/2$. Let $v_{d}$ denote the volume of a unit ball and define $C_{d}:=2^{d}/v_{d}$ then
$N_{[]}(uM^{(n)}_{u},\mathcal{F}(\Omega_{n}))\leq N(u,\Omega)\leq D(u,\Omega)\leq C_{d}\text{Vol}(\tilde{\Omega}_{n})u^{-d}.$
@@
Due to the restriction that $n\phi_{n}^{2}\rightarrow\infty$, we have $\phi_{n}^{-1}=o(n^{1/2})$, thus every term in the right-hand side of (26) are of order $\log n$, apart from $\log M_{u}^{(n)}$. By assumption, there exists $q\in(0,1)$ such that $\forall u\in(0,\phi_{0})$,
@@
for $n$ sufficiently large which implies (8).∎

### v3

@@ 5 Examples
We have seen in the previous section that empirically the approximate $\hat{\epsilon}_{n}$ follows the decreasing trend as in Theorem 4. In this section, we give a few examples where Theorem 4 is applicable theoretically. The verification of Assumptions 1 to 3 across different models shares a few common procedures. For Assumption 4 to hold for regression models (with covariates), one needs to control the leverage of the covariates $\bm{W}$. It suffices to assume $\|\bm{W}\|_{\infty}\leq n^{\iota}$, for some $\iota\in(0,1)$ that depends on the choice of $\phi_{n}$. However, for simplicity, we will still assume $\|\bm{W}\|_{\infty}\leq 1$, as we did in the simulations in Section 4. With $\|\bm{W}\|_{\infty}\leq 1$, Assumption 4 can be trivially satisfied by correctly choosing a heavy-tailed contamination density $g$, and Assumption 5 is usually easy to check for linear models.
Due to space constraints, the detailed analysis is deferred to Appendix C. Typically, we can set $\phi_{n}$ and $\psi_{n}$ in Proposition 3 to decay at the same rate, with $\phi_{n},\psi_{n}\propto n^{-1/q}$ and $p_{n}=n^{-1/2q}$, for some $q>4$, which gives
$\epsilon_{n}\sim n^{-1/2q},\qquad\delta_{n}\sim\exp(-\tfrac{1}{2}c_{1}n^{1-2/q}),$
where $c_{1}$ is a fixed constant.
@@ 6 Discussion
Differential privacy mechanisms in statistical inference can be roughly divided into two categories: 1. Local models, where data are privatised before collection and inference is conducted on the perturbed dataset; 2. Central models, where data are collected and processed without noise, but the result is privatised before release.
@@
The ”local model”, or ”local differential privacy” as defined in Erlingsson et al. (2014), quantifies the level of differential privacy provided by the mechanism on ”datasets” with only one data point. Thus, contamination can be done individually before the dataset is gathered. Huber’s contamination model is a popular approach to local differential privacy. Our work, however, considers differential privacy from a central model perspective despite using a local privatisation mechanism. Notably, the local differential privacy level only depends on the contamination probability $p_{n}$, but $p_{n}\rightarrow 0$ as $n\rightarrow\infty$. From the local model perspective, our setup will be asymptotically non-private as $n\rightarrow\infty$. However, from the central model perspective, with a trusted curator, our approach can ensure full differential privacy almost surely with an asymptotically zero contamination rate.
@@
Huber’s contamination model for robust estimation by contaminating the dataset with a heavy-tailed distribution also introduces differential privacy to posterior sampling, similar to central models that directly inject Laplace/Gaussian noise into the output. By contaminating the dataset with a heavy-tailed distribution, we can lift the assumption on bounded observation space. We further extend the result, through a probabilistic argument, to lift the bound on the parameter space and upper bound the rate of convergence for $\epsilon$ and $\delta$.
@@
Through simulation, a practitioner can estimate the level of differential privacy without having to hand-compute all the constants in the theorem. Despite having several non-trivial assumptions for the theorem, we show in Section 5 and Appendix C how the assumptions could be verified mathematically and how the convergence rate of $\epsilon$ and $\delta$ can be derived under a few common regression setups.
The analysis presented in this work relies on the assumption that i.i.d. samples are drawn exactly from the target distribution. The bounds may be extended to non-exact samplers by applying the results of Minami et al. (2016), provided that the algorithm admits a provable convergence guarantee. The study of convergence properties of Markov chain Monte Carlo methods is an active area of research (e.g., Durmus and Moulines (2017); Andrieu et al. (2022, 2025)). Integrating such convergence results into our framework represents a promising avenue for future work.
Our main result (Theorem 4) holds uniformly for all dataset sizes $n$, provided the underlying assumptions are satisfied. Nevertheless, this should be interpreted as an asymptotic result since $n$ has to be large for some assumptions to hold in general. This may be improved by using the results of Ghosal and Van der Vaart (2017), for which we include a detailed discussion in Appendix D.1.
Furthermore, posterior contraction results are also established in broader Bayesian inference frameworks, including misspecified models and infinite-dimensional parameter spaces. While the specific assumptions employed here may not directly apply in those contexts, analogous differential privacy bounds can be derived under appropriately modified assumptions, provided a suitable contamination density can be constructed. We defer such generalisations to future work.
@@
$\mathbb{P}^{*}_{\theta^{*}}\left(\sup_{\theta\in A_{n}^{c}\cap\Omega_{n}}\prod_{i=1}^{n}d(\bm{x}_{i},\theta)\geq\exp({-c_{1}n\phi_{n}^{2}})\right)\leq(1+L(\phi_{n}))(1+N(\phi_{n}))\exp({-c_{2}n\phi_{n}^{2}}),$
@@
If in addition $\exp\left(-c_{2}n\phi_{n}^{2}\right)\leq\frac{1}{2},$ then
@@
Setting $p^{\prime}=0$, we have
@@
See Lemma 2 of Shen and Wasserman (2001); the proof holds for any fixed $p\in[0,1)$. ∎
@@
For $\alpha\in(0,1]$, $S_{p,\alpha}(t)=\{\theta\in\Omega:\rho_{p,\alpha}(\theta^{*},\theta)\leq t\}$, then $S_{0,\alpha}(t)\subseteq S_{p,\alpha}(t)$.
@@
Then applying Lemma 10 with the above inequality gives the bound. ∎
@@
By Lemmas 12 and 13, and recall that $t_{n}=c_{1}\phi_{n}^{2}/8$ defined in Assumption 2,
@@
and from Lemma 14 that $\eta_{n}\leq e^{2(1-p_{n})T_{n}\psi_{n}/p_{n}}.$ Let $\delta_{n}=\tfrac{4}{c_{3}}e^{-\frac{1}{2}c_{1}n\phi_{n}^{2}}$, $\varepsilon_{n}=2(1-p_{n})T_{n}\psi_{n}/p_{n}$ and $\epsilon_{n}=\varepsilon_{n}+\log(\delta_{n}+\exp(\varepsilon_{n})).$ Note that since $\bm{X}$ and $\bm{Z}$ have the same cardinality, applying Lemma 13 again to $m_{n}(A_{n}^{c},\bm{Z})$, we have
$\mathbb{P}\Bigl(\mathbb{P}_{\pi_{n}}(S|\bm{X})\leq e^{\epsilon_{n}}\mathbb{P}_{\pi_{n}}(S|\bm{Z})+\delta_{n}\Bigr)\geq$  $\,1-2\mathbb{P}\left(\frac{m_{n}(A_{n}^{c},\bm{X})}{m_{n}(\Omega,\bm{X})}\geq\frac{4}{c_{3}}e^{-\frac{1}{2}c_{1}n\phi_{n}^{2}}\right)$
$\geq$  $\,1-4e^{-nc_{1}\alpha\phi_{n}^{2}/8}-2C(n,\phi_{n})e^{-c_{2}n\phi_{n}^{2}}-2c_{4}e^{-c_{1}n\phi_{n}^{2}}$
$\geq$  $\,1-C^{\prime}(n,\phi_{n})e^{-nc_{5}\phi_{n}^{2}},$
where $c_{5}=\min\{c_{1}\alpha/8,c_{2}\}$ and $C^{\prime}(n,\phi_{n})=4+2C(n,\phi_{n})+2c_{4}$. Note that by definition of $C(n,\phi_{n})$ from Lemma 13, $C^{\prime}(n,\phi_{n})=4+40(\log_{2}n)^{2}+2c_{4}$. On the other hand, if $n$ is large enough, then $C^{\prime}(n,\phi_{n})=12+2c_{4}$ is independent of $n$, and the proof is complete. ∎
@@
Let $H(\bm{x}):=\left[\nabla_{\theta}f(\bm{x};\theta^{*})\right]\left[\nabla_{\theta}f(\bm{x};\theta^{*})\right]^{\top}$ and let $H(\bm{x})]_{i,j}$ denote its $(i,j)$-th entry. Consider the sets $A_{i,j}^{+}:=\{x:[H(\bm{x})]_{i,j}>0\}$ and similarly $A_{i,j}^{-}:=\{x:[H(\bm{x})]_{i,j}<0\}$. Then we have
$0<\left|\frac{[H(\bm{x})]_{i,j}}{k_{p_{n}}(\bm{x};\theta^{*})f(\bm{x};\theta^{*})}\right|\leq\frac{[H(\bm{x})]_{i,j}}{(1-p_{n})f^{2}(\bm{x};\theta^{*})},\text{ for }x\in A_{i,j}^{+},$
@@
$0<\left|\frac{[H(\bm{x})]_{i,j}}{k_{p_{n}}(\bm{x};\theta^{*})f(\bm{x};\theta^{*})}\right|\leq-\frac{[H(\bm{x})]_{i,j}}{(1-p_{n})f^{2}(\bm{x};\theta^{*})},\text{ for }x\in A_{i,j}^{-}.$
Note that $A_{i,j}^{+}$ and $A_{i,j}^{-}$ are both measurable with
$I(\theta^{*})]_{i,j}=\int_{A_{i,j}^{+}\cup A_{i,j}^{-}}\frac{[H(\bm{x})]_{i,j}}{f^{2}(\bm{x};\theta^{*})}f(\bm{x};\theta^{*})\mathtt{d}x.$
Thus, applying the dominated convergence theorem on both $A_{i,j}^{+}$ and $A_{i,j}^{-}$,
$\lim_{n\rightarrow\infty}\left[I_{p_{n}}(\theta^{*})\right]_{i,j}$ $\,=\lim_{n\rightarrow\infty}(1-p_{n})^{2}\left(\int_{A_{i,j}^{+}}\frac{[H(\bm{x})]_{i,j}}{k_{p_{n}}(\bm{x};\theta^{*})f(\bm{x};\theta^{*})}f(\bm{x};\theta^{*})\mathtt{d}x\right$.
$\,\qquad\qquad\qquad\qquad\left.+\int_{A_{i,j}^{-}}\frac{[H(\bm{x})]_{i,j}}{k_{p_{n}}(\bm{x};\theta^{*})f(\bm{x};\theta^{*})}f(\bm{x};\theta^{*})\mathtt{d}x\right)$
$\,=\int_{A_{i,j}^{+}\cup A_{i,j}^{-}}\frac{[H(\bm{x})]_{i,j}}{f^{2}(\bm{x};\theta^{*})}f(\bm{x};\theta^{*})\mathtt{d}x=[I(\theta^{*})]_{i,j}.$
@@ B.1 Results in Main Text
[…]
@@ B.1.1 Proof for Lemma 2
The first half of this proof is based on several results commonly seen in the empirical processes literature (with some variation), see, for instance, Section 2.8 of Van der Vaart and Wellner (1996).
@@
$N_{[]}(uM^{(n)}_{u},\mathcal{F}(\Omega_{n}))\leq C_{d}\text{Vol}(\tilde{\Omega}_{n})u^{-d},$ (22)
@@
Let $\{\theta_{1},\dots,\theta_{k}\}$ be a $u$-cover of $\Omega_{n}$ with respect to $\|\cdot\|_{\infty}$. Consider the following functions, for $i=1,\dots,k$,
@@
Thus the set of $L_{i},U_{i}]$ forms a $uM^{(n)}_{u}-$bracketing for $\mathcal{F}(\Omega_{n})$ under the Hellinger distance $h(\cdot,\cdot)$, hence,
@@
To complete (22), we recall that $N(u,\Omega_{n})\leq D(u,\Omega_{n})$ always holds. If $\{\theta_{1},\dots,\theta_{k}\}$, $k:=D(u,\Omega_{n})$, is a set of $u$-separated points, then the balls of radius $u/2$ centred at $\{\theta_{1},\dots,\theta_{k}\}$ are disjoint and contained in $\tilde{\Omega}_{n}:=\{\theta\in\Omega:\exists\theta^{\prime}\in\Omega_{n},\|\theta-\theta^{\prime}\|_{2}\leq u/2\}$. Thus $D(u,\Omega_{n})$ is upper bounded by the ratio between the volume of the space $\tilde{\Omega}_{n}$ and the volume of a ball of radius $u/2$. Let $v_{d}$ denote the volume of a unit ball, and define $C_{d}:=2^{d}/v_{d}$, then
$N_{[]}(uM^{(n)}_{u},\mathcal{F}(\Omega_{n}))\leq N(u,\Omega_{n})\leq D(u,\Omega_{n})\leq C_{d}\text{Vol}(\tilde{\Omega}_{n})u^{-d}.$
@@
Due to the restriction that $n\phi_{n}^{2}\rightarrow\infty$, we have $\phi_{n}^{-1}=o(n^{1/2})$, thus every term on the right-hand side of (23) are of order $\log n$, apart from $\log M_{u}^{(n)}$. By assumption, there exists $q\in(0,1)$ such that $\forall u\in(0,\phi_{0})$,
@@
for $n$ sufficiently large, which implies (8).∎